What mass of water is produced from the complete combustion of 4.00×10−3 g of methane?
1
Expert's answer
2016-09-17T09:01:13-0400
Methane combustion follows the equation CH4 + 2O2 = CO2 + 2H2O. Thus, 4.00*10-3 g of methane is 4.00*10-3 g/16 g/mol = 2.5*10-4 mol and it produces 5*10-4 mol of water or 5*10-4 mol * 18 g/mol = 9*10-3 g of it.